A carpenter has a board that is 15 ft long. He needs to cut it into pieces so that one piece is 3 ft shorter than the other piece. Find the length of each piece.
Found 2 solutions by Boreal, FrankM:
Answer by Boreal(15235)

longer piece is x
shorter piece is x-3
they add to 15
x+(x-3)=15
2x-3=15
2x=18
x=9 ft.
x-3=6 ft.

Answer by FrankM(1040)

Using algebra (of course),
X is one, X+3 is the other.
X+X+3=15
2X+3=15
X=6 and X+3=9